# Growth driven phase transitions in Zinc Oxide nanoparticles through machine-learning assisted simulations

### v1

@@ II.2 Growth Simulations
For the subsequent deposition, a spherical shell with a radius of 25Å is defined around the nanoparticle, within which a zinc atom and an oxygen atom are introduced at fixed time intervals and at random positions. The newly introduced atoms are assigned initial velocities directed toward the center of mass of the system in order to enforce deposition and growth. Atoms are deposited with rate of one pair of Zinc and Oxygen atoms every picosecond which is by far slower than with ab initio MD while remaining several orders of magnitudes faster than experiments Zhao2016Apr . Between two depositions, a molecular dynamics (MD) simulation of the system is carried out at a given temperature in the NVT ensemble. We note that the stochastic nature of the process leads to some depositions being unsuccessful ie. the deposited atoms bounce on the surface and return in the gas phase. Those particular atoms are removed from the simulation box and the the growth simulation is terminated when 200 ZnO depositions were attempted. For each initial cluster and thermodynamic condition, 10 different growth simulations are performed starting from different initial conditions for the velocities and the random positions of the deposited atoms. The trajectories are generated using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS).Thompson2022Feb All the simulations are conducted with a timestep of 1 fs in the NVT ensemble with a temperature controlled by the Nosé–Hoover thermostat using a damping parameter of 100 fs.
@@ III.2 General behavior
To assess whether the nanoparticle (NP) size is the key parameter governing the BCT-to-WRZ structural transition, the NP size at the transition point has been plotted as a function of the initial BCT seed size in Fig. 5, with data averaged over 10 independent simulation simulations. This transition point is defined as the moment when the number of atoms in the WRZ-like configuration surpasses those in the BCT-like configuration. As previously noted, temperature appears to have a minimal influence on the transition behavior. In contrast, the crystalline cluster size at the transition tends to increase with the initial seed size, suggesting a correlation between the two. However, this observation should be interpreted with caution, as the large error bars indicate significant variability in the transition size across different simulations. Therefore, while the overall NP size may contribute to triggering the transition, it is unlikely to be the sole determining factor.
Figure 5: Size of the NP in number of atoms at the BCT-to-WRZ transition as a function of the initial BCT seed at 500K, 700K and 900K, averaged over the 10 independent simulations.
@@ III.3 Microscopic picture of the growth mediated phase transition
Figure 6: Selected snapshots of the nanoparticle during the deposition simulations. (a) to (e): WRZ-472 initial seed. (f) and (j): BCT-648 initial seed. In each case, the top images show the entire NP with the deposited atoms in different color codes. The bottom images show the crystalline cores with the BCT atoms in blue and the WRZ atoms in green.
@@
To gain deeper insight into the mechanisms governing the BCT-to-WRZ transition, it is instructive to first recall the distinct nature of their terminal facets. The opposing (100) end facets of BCT-type nanoparticles are stoichiometric and therefore non-polar. In contrast, the (0001) and (000–1) end facets of WRZ-type nanoparticles are polar and thus require an excess of O and Zn atoms, respectively, to achieve electrostatic stabilization.Goniakowski2007Dec Consequently, the BCT-to-WRZ transformation inherently involves the development of ionic separation, which manifests as opposite compensating ionic excesses at the opposing NP’s facets. To monitor this effect during deposition simulations, the Zn excess was systematically quantified at both terminations along the principal (long) axis of the NPs. Figures 7 and 8 illustrate representative examples of this analysis, displaying the temporal evolution of the number of atoms associated with each crystalline phase within the nucleus, along with the corresponding variation of Zn excess for WRZ and BCT initial seeds, respectively.
@@
Focusing first on the growth of nanoparticles from WRZ seeds — where, as described in the Methods section, the required ionic excess was incorporated into the initial configurations — we find that particles initiated from the larger seeds (Figs. 7(b, c)) exhibit steady growth of their WRZ crystalline cores at comparable rates ($\Delta N_{crys}/\Delta N_{deposited}\sim$ 0.25). In both cases, the initial Zn excess is largely preserved throughout deposition, indicating that the added atoms primarily contribute to elongation of the crystalline core and development of a less ordered peripheral zone, with only minor impact on core width. This preservation of Zn excess also suggests the absence of additional polarity effects induced by the small particle size GoniakowskiPRB2011 ; NogueraCR2013 . In contrast, in smaller seeds (Figs. 7(a, d)), the initial WRZ phase disappears nearly completely due to its thermodynamic instability at the smallest particle sizes. Its disappearance is followed by a rapid decrease of the initial Zn excess, thus demonstrating that ionic mobility during deposition efficiently modulates local NP composition in response to changes in the character of the exposed facets. Partial recovery of the WRZ phase is observed at later growth stages (Fig. 7(a)), although this recovery can be significantly delayed by the presence of a well-developed BCT core (Fig. 7(d)). In both these cases, the overall WRZ fraction remains low, resulting in only marginal ionic separation.
@@
Figure 8: ((a) to (d): Examples of the evolution of the number of atoms of the BCT (blue) and WRZ (green) type inside the crystalline nucleus as a function of the number of deposited atoms for different BCT initial seeds at 700 K. (e) to (h): Excess of Zn atoms at opposite ends of the nanoparticle.
@@
Turning to nanoparticle growth from BCT seeds (Fig. 8), we find that the progressive emergence of the WRZ phase — observed consistently across all cases — correlates with markedly different behaviors in Zn excess. In cases 8(a and c), where the BCT-to-WRZ transformation occurs at an early stage of deposition, the steady growth of the WRZ crystalline core closely matches that of larger WRZ-seeded particles (Figs. 7(b and c)), both in growth rate ($\Delta N_{crys}/\Delta N_{deposited}\sim$ 0.25) and in the complete absence of residual BCT content. These systems also exhibit rapid establishment of stable ionic separation, with Zn excess and deficiency localized at opposite nanoparticle ends, consistent with the polarity of the two basal facets of the growing WRZ core. In contrast, in case 8(b), the formation and growth of the WRZ core are substantially delayed by the persistence of well-defined BCT grains. Steady WRZ growth commences only near the end of the simulation and is accompanied by a sharp onset of ionic separation. This behavior closely resembles that of small WRZ-seeded nanoparticles (Figs. 7(a and d)), confirming that for the smallest seeds their initial structural configuration may be of limited importance since the high structural flexibility at the considered temperature allows thorough reorganization during growth. Finally, case 8(d) presents a less straightforward scenario. Here, competition between the initial BCT phase and the emerging WRZ phase results in a delayed and overall less steady increase of WRZ content, with residual BCT persisting throughout the entire simulation — unlike in cases 8(a, c) and 7(b, c). Most surprisingly, despite the WRZ phase ultimately becoming largely dominant, no detectable ionic separation is observed. This apparent discrepancy is due to the orientation of the crystalline WRZ core which forms in this nanoparticle. Indeed, we observed that its polar direction is perpendicular to the long particle axis and the present Zn excess analysis becomes inconclusive in this case.

### v2

@@ II.1 Charge included Physical Lassolars Interaction Potential (PLIP+Q)
The PLIP+Q model was previously tested for many physical properties including lattice spacing, polymorph stability, radial distribution functions, surface energies and nanoparticle stability Salazar2024Aug . In addition, to specifically verify our usage in growth conditions, we performed additional single point calculations [See SI A].
@@ II.2 Growth Simulations
For the subsequent deposition, a spherical shell with a radius of 25Å is defined around the nanoparticle, within which a zinc atom and an oxygen atom are introduced at fixed time intervals and at random positions. The newly introduced atoms are assigned initial velocities directed toward the center of mass of the system in order to enforce deposition and growth. Atoms are deposited with rate of one pair of Zinc and Oxygen atoms every 10 picoseconds which is by far slower than with ab initio MD while remaining several orders of magnitudes faster than experiments Zhao2016Apr . It is to be noted that the chosen deposition rate is faster than what can typically be achieved with classical force fields. This is due to the computational cost of the PLIP+Q model, which requires 20 hours to produce 1 ns of simulation time using 32 CPUs under our growth conditions. Meanwhile, we also tested deposition rate that were five times slower and obtained qualitatively similar results [See SI B]. Between two depositions, a molecular dynamics (MD) simulation of the system is carried out at a given temperature in the NVT ensemble. We note that the stochastic nature of the process leads to some depositions being unsuccessful ie. the deposited atoms bounce on the surface and return in the gas phase. Those particular atoms are removed from the simulation box and the the growth simulation is terminated when 200 ZnO depositions were attempted. For each initial cluster and thermodynamic condition, 10 different growth simulations are performed starting from different initial conditions for the velocities and the random positions of the deposited atoms. In addition, we also measured the surface mobility of the atoms and verified that our conditions lead to enough mobility for atomic reconstruction [See SI C]. The trajectories are generated using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS).Thompson2022Feb All the simulations are conducted with a timestep of 1 fs in the NVT ensemble with a temperature controlled by the Nosé–Hoover thermostat using a damping parameter of 100 fs.
@@ III.2 General behavior
To assess whether the nanoparticle (NP) size is the key parameter governing the BCT-to-WRZ structural transition, the number of deposited atoms at the transition point has been plotted as a function of the initial BCT seed size in Fig. 5, with data averaged over 10 independent simulation simulations. This transition point is defined as the last moment when the number of atoms in the BCT-like configuration surpasses those in the WRZ-like configuration. As previously noted, temperature appears to have a minimal influence on the transition behavior. More surprisingly, it seems that the number of deposited atoms at the transition point does not really depends on the size of the seed as well. We note that the measured error bars are relatively large. But, using a convergence of the mean analysis, we can show that 10 independent simulations is enough [see SI D] which means that the large error bars are only due to the stochastic nature of the crystallization process.
Figure 5: Number of deposited atoms at the BCT-to-WRZ transition as a function of the initial BCT seed at 500K, 700K and 900K, averaged over the 10 independent simulations.
Before detailing the microscopic process occurring during the growth, we would like to return to the PLIP+Q usage. In particular, a primary advantage of PLIP+Q over conventional MLIP models is its explicit decomposition of interactions into 2-body, 3-body, N-body, and electrostatic components. We leveraged this framework to gain physical insights into the nature of atomic interactions in zinc oxide. By analyzing our growth simulations, we calculated the contribution of each term — averaging the atomic energy and forces across all atoms in all of the 75 NP configurations considered during the validation of the PLIP+Q models done in SI A (see Fig. 6). Our analysis revealed several key findings. Firstly, not all components contribute with the same sign to the total interaction. Terms are denoted as positive when they align with the sign of the total interaction and negative when they oppose it. Secondly, energy and forces are not partitioned equivalently. Some interaction terms contribute significantly to the total energy while remaining near a local minimum for the corresponding gradients, resulting in minimal force contributions. Thirdly, comparing the four terms shows that Coulombic interactions are generally the most dominant in the energy landscape. Yet, when summing homonuclear and heteronuclear contributions, the Coulombic forces largely cancel out [See Fig 6.(b) ”All” contributions]. Consequently, 2-body and N-body terms become the primary drivers of the total force. This cancellation likely stems from the use of uniform electrostatic charges for the Zn and O atoms.
Figure 6: Contributions of 2-body, 3-body, N-body, and electrostatic terms to the interaction energies and interatomic forces for oxygen–oxygen, zinc–zinc, and oxygen–zinc pairs, obtained by averaging over all atoms and the entire growth simulation time.
@@ III.3 Microscopic picture of the growth mediated phase transition
Figure 7: Selected snapshots of the nanoparticle during the deposition simulations. (a) to (e): WRZ-472 initial seed. (f) and (j): BCT-648 initial seed. In each case, the top images show the entire NP with the deposited atoms in different color codes. The bottom images show the crystalline cores with the BCT atoms in blue and the WRZ atoms in green. In Fig. 7(a,f), the dashed lines correspond to the planes used for computing the Zn excess.
@@
To gain deeper insight into the mechanisms governing the BCT-to-WRZ transition, it is instructive to first recall the distinct nature of their terminal facets. The opposing (100) end facets of BCT-type nanoparticles are stoichiometric and therefore non-polar. In contrast, the (0001) and (000–1) end facets of WRZ-type nanoparticles are polar and thus require an excess of O and Zn atoms, respectively, to achieve electrostatic stabilization.Goniakowski2007Dec Consequently, the BCT-to-WRZ transformation inherently involves the development of ionic separation, which manifests as opposite compensating ionic excesses at the opposing NP’s facets. To monitor this effect during deposition simulations, the charge imbalance was systematically quantified at both terminations. For that purpose, we first remove any stochastic translation and rotation of the system. We then split the NPs into two equal halves by a plane passing through the center of the nanoparticle and oriented parallel to the original basal facets [See Fig. 7(a,f)]. For each half, we calculate the difference between the number of zinc and oxygen atoms, which we term Zn excess. In few of the BCT cases, growth can also occur perpendicularly which we also considered by using a perpendicular plane [See Fig. 7(f)]. Figures 8 and 9 illustrate representative examples of this analysis, displaying the temporal evolution of the number of atoms associated with each crystalline phase within the nucleus, along with the corresponding variation of Zn excess for WRZ and BCT initial seeds, respectively.
@@
Focusing first on the growth of nanoparticles from WRZ seeds — where, as described in the Methods section, the required ionic excess was incorporated into the initial configurations — we find that particles initiated from the larger seeds (Figs. 8(b, c)) exhibit steady growth of their WRZ crystalline cores at comparable rates ($\Delta N_{crys}/\Delta N_{deposited}\sim$ 0.25). In both cases, the initial Zn excess is largely preserved throughout deposition, indicating that the added atoms primarily contribute to elongation of the crystalline core and development of a less ordered peripheral zone, with only minor impact on core width. This preservation of Zn excess also suggests the absence of additional polarity effects induced by the small particle size GoniakowskiPRB2011 ; NogueraCR2013 . In contrast, in smaller seeds (Figs. 8(a, d)), the initial WRZ phase disappears nearly completely due to its thermodynamic instability at the smallest particle sizes. Its disappearance is followed by a rapid decrease of the initial Zn excess, thus demonstrating that ionic mobility during deposition efficiently modulates local NP composition in response to changes in the character of the exposed facets. Partial recovery of the WRZ phase is observed at later growth stages (Fig. 8(a)), although this recovery can be significantly delayed by the presence of a well-developed BCT core (Fig. 8(d)). In both these cases, the overall WRZ fraction remains low, resulting in only marginal ionic separation. We note that the process shown in Fig. 7(a) and Fig. 7(d) occurs only five times across our 90 simulations. We observe that, in all cases where growth of WRZ phase is substantially delayed, the Zn excess is not being maintained during the atom-by-atom deposition [See SI E]. In contrast, as we will show next, emergence of a Zn excess can actually induce a phase transition when starting from BCT seeds.
@@
Figure 9: ((a) to (d): Examples of the evolution of the number of atoms of the BCT (blue) and WRZ (green) type inside the crystalline nucleus as a function of the number of deposited atoms for different BCT initial seeds at 700 K. (e) to (h): Excess of Zn atoms at opposite ends of the nanoparticle. In dashed lines, the Zn excess in the perpendicular direction.
@@
Turning to nanoparticle growth from BCT seeds (Fig. 9), we find that the progressive emergence of the WRZ phase — observed consistently across all cases — correlates with markedly different behaviors in Zn excess.In cases 9(a and c), where the BCT-to-WRZ transformation occurs at an early stage of deposition, the steady growth of the WRZ crystalline core closely matches that of larger WRZ-seeded particles (Figs. 8(b and c)), both in growth rate ($\Delta N_{crys}/\Delta N_{deposited}\sim$ 0.25) and in the complete absence of residual BCT content. These systems also exhibit rapid establishment of stable ionic separation, with Zn excess and deficiency localized at opposite nanoparticle ends, consistent with the polarity of the two basal facets of the growing WRZ core. In contrast, in case 9(b), the formation and growth of the WRZ core are substantially delayed by the persistence of well-defined BCT grains. Steady WRZ growth commences only near the end of the simulation and is accompanied by a sharp onset of ionic separation. This behavior closely resembles that of small WRZ-seeded nanoparticles (Figs. 8(a and d)), confirming that for the smallest seeds their initial structural configuration may be of limited importance since the high structural flexibility at the considered temperature allows thorough reorganization during growth. Finally, case 9(d) presents a less straightforward scenario. Here, competition between the initial BCT phase and the emerging WRZ phase results in a delayed and overall less steady increase of WRZ content, with residual BCT persisting throughout the entire simulation. In this case, the polar direction of the WRZ phase still emerges but in the perpendicular direction with respect to the initial NP long axis [See dashed lines in Fig 9(h)].
[…]